The following data is given for X and Y.
X : 2 3 5 7 10 12
Y: 5 9 12 14 15 20 calculate Karl Pearson’s coefficient of correlation
between X &Y
1
Expert's answer
2020-08-18T16:14:55-0400

Karl Pearson's correlation coefficient (r) formula is given by -


"r = \frac{\sum(X-\bar{X})(Y-\bar{Y})}{\sqrt{(\sum(X-\bar{X})^2)\ *\ ({\sum(Y-\bar{Y})^2)}}}"